On a map, 3/4 inch represents 30 miles. If two are 4 inches apart on the map, how many miles apart are they?
Answer by jim_thompson5910(35256)

(3/4)/30 = 4/x

(3/4)x = 30*4

(3/4)x = 120

3x = 120*4

3x = 480

x = 480/3

x = 160

The two towns are 160 miles apart.